OpenStudy (anonymous):

5x+15 over 8 divided by 4x+12 over 12

OpenStudy (anonymous):

\[\frac{\frac{1}{8} (15+5 x)}{\frac{1}{12} (12+4 x)}=\frac{\frac{1}{8} 5 (3+x)}{\frac{1}{12} 4 (3+x)}=\frac{\frac{5}{8}}{\frac{1}{3}}=\frac{15}{8} \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np Thank you for the medal.